# Embedded Hyperspectral Band Selection with Adaptive Optimization for Image Semantic Segmentation

### v1

@@ 4.4 Experimental Design and Evaluation
In order to test our method and the benchmark methods in a robust and accurate way, we ran each experiment with 10-fold cross-validation. In addition, as most of the baseline methods are not embedded, we first run the various band selection models as a first stage and then used the selected bands from each method to train the downstream task model via the baseline semantic segmentation CNN model. This ensured a proper apples-to-apples comparison of the various band selection methods.
For evaluation, we compared the accuracy of the various methods for various target number of bands. In addition, in order to compare methods across the whole range of numbers of selected bands, we introduced an new evaluation metric which calculated the area under the bands-performance curve. This metric was inspired by the common Area Under the Curve (AUC) metric used to evaluate the performance of a binary classification model, particularly in the context of a Receiver Operating Characteristic (ROC) curve. The AUC provides a single scalar value that summarizes the performance of a classification model across various classification thresholds and is widely used metric for evaluating and comparing binary classification models. Similarly, in our setting, the AUC provides a single scalar value that summarizes the performance of a band selection model across various number of bands selected allowing us to compare different models by a single value.
For testing our EHBS model, in each experiment we run the model multiple times with different regularization factor ($\lambda$) values in the range 0.2 to 2.4 applying a simple heuristic search till the desired target number of selected bands was met. The number of bands that were used and the corresponding accuracy results were saved for each run. Overall accuracy for a given target number of bands was calculated by averaging the accuracy across all folds.
@@ 5 Results
Figure 6: Illustration of the selected bands for the Salinas dataset across various target number of bands. Each horizontal ”row” represents selection of bands in a single run of EHBS with the selected bands marked in black. The top row corresponds to the selection of 8 bands and the bottom row to a run with 48 bands selected.
@@ Methods
EHBS Waludi Walumi ISSC
BS-NETS
Conv
BS-NETS
FC
@@
EHBS Waludi Walumi ISSC
BS-NETS
Conv
BS-NETS
FC
@@
EHBS Waludi Walumi ISSC
BS-NETS
Conv
BS-NETS
FC
@@
EHBS Waludi Walumi ISSC
BS-NETS
Conv
BS-NETS
FC
@@
Figure 7 shows a graph comparing the accuracy of our proposed model, EHBS, as well as the various baseline band selection models on the PaviaU dataset with a patch size of 7x7 over multiple target number of bands. As can be seen from the graph our proposed model outperforms the baseline models in most of the data points and is comparable to the best baseline model in the others. it also shows the corresponding AUC metric for the various models. Overall, our EHBS model achieved the highest AUC value surpassing the performance of all other models. Figure 8 shows the accuracy results of our proposed method over the whole range of target band selection for the PaviaU dataset and a 7x7 patch size. For this specific dataset, close to optimal performance is obtained by using the 20 best-selected bands (or $\sim$20of the bands). Performance is stable with only slight improvement from this point till the use of the full hyperspectral input.
Table 1 shows a comparison table including accuracy results and AUC results over the PaviaU dataset over two different patch sizes (7x7 and 11x11) and selected target number of bands. As baseline models we chose the top performing models on the PaviaU 7x7 experiments. Table 2 shows a similar comparison table for the Salinas dataset. For all datasets and patch sizes, EHBS obtained the highest AUC score and obtained the best accuracy for all Salinas experiments and for the vast majority of PaviaU experiments. In the few PaviaU experiments for which EHBS was not the top performing method it obtained comparable accuracy results to the top performing method.
Figure 6 illustrates the selected bands for the Salinas dataset across various target number of bands. EHBS systematically ignores areas with low contribution for the downstream task performance and the prominent band selection areas remain consistent across different runs.
@@ 6 Discussion and Conclusion
Following, we highlight several key findings and implications derived from our experiments. Firstly, our proposed method demonstrates clear superiority across two diverse datasets and various experimental settings, underscoring its robustness and effectiveness. We conducted comparisons among the methods in various settings, including different patch sizes, datasets, and the number of selected bands. We observed that for larger patch sizes, our method outperformed others by an even wider margin. We attribute this phenomenon to the presence of more contextual information and a higher model complexity. Consequently, employing an embedded model enables the model to discover the optimal bands that effectively capture the intricate non-linear interactions among the features. Additionally, our findings suggest that our model tends to outperform baseline models when confronted with an abundance of training data and a higher number of bands to select. We thus hypothesize that our method would perform even better on larger datasets, with larger patch sizes, and with different architectures that hold a global view per band. However, exploration of these scenarios will be reserved for future work. We demonstrated the strength of our model while integrated into CNN-based deep learning model for hyperspectral semantic segmentation. The seamless integration of our method into an existing deep learning model that was achieved by simply adding a selection layer after the input enhances its practicality and ease of use. This simplicity of implementation coupled with the robust performance positions our method as a promising and practical solution for various machine learning and computer vision tasks.
@@
These results also demonstrate the robustness of our proposed method along multiple settings and various target number of bands. The method performs well and consistently without the need to tweak or modify the model regardless of the target number of bands - be it only a few number of bands or when aiming for a large number of selected bands. This also allows researchers to get a good understanding of the performance curve and choose the desired tradeoff given the application setting.
@@
After our initial experimentation, we further tested more optimizers applying an extensive search for an optimal learning rate. As can be seen in Table 3, when using the Adam optimer the results are sensitive to the learning rate and finding the right learning rate is crucial. It turns out that on the PaviaU dataset with a 7x7 path, EHBS with the Adam optimizer can achieve even better performance than EHBS with DoG. However, this is a retrospective result obtained over our testset and not via tuning on an held out evaluation set. Given that the results achieved with the DoG optimizer are comparable to the optimal outcomes obtained using the Adam optimizer, we advocate for the adoption of DoG. Notably, DoG eliminates the need for parameter tuning, ensuring a more robust and reliable performance.
@@ 7 Acknowledgement
This work was partially supported by the Chief Scientist of the Israeli Ministry of Agriculture grant number12-03-0010. We would also like to thank Nadav Spitzer for his help in implementing the baseline feature selection models and to Ofir Lindenbaum and Jacob Goldberger for providing input and feedback on our work.

### v2

@@ 4.4 Experimental Design and Evaluation
In order to test our method and the benchmark methods robustly and accurately, we ran each experiment with 10-fold cross-validation. In addition, as most of the baseline methods are not embedded, we first ran the various band selection models as a first stage and then used the selected bands from each method to train the downstream task model via the baseline semantic segmentation CNN model. This ensured a proper apples-to-apples comparison of the various band selection methods.
For evaluation, we compared the accuracy of the various methods for various target numbers of bands. In addition, in order to compare methods across the whole range of numbers of selected bands, we introduced a new evaluation metric that calculated the area under the bands-performance curve. This metric was inspired by the common Area Under the Curve (AUC) metric used to evaluate the performance of a binary classification model, particularly in the context of a Receiver Operating Characteristic (ROC) curve. The AUC provides a single scalar value that summarizes the performance of a classification model across various classification thresholds and is a widely used metric for evaluating and comparing binary classification models. Similarly, in our setting, the AUC provides a single scalar value that summarizes the performance of a band selection model across various numbers of bands selected, allowing us to compare different models by a single value.
For testing our EHBS model, in each experiment, we ran the model multiple times with different regularization factor ($\lambda$) values in the range 0.2 to 2.4, applying a simple heuristic search till the desired target number of selected bands was met. The number of bands that were used and the corresponding accuracy results were saved for each run. Overall accuracy for a given target number of bands was calculated by averaging the accuracy across all folds.
@@ 5 Results
Figure 6: Illustration of the selected bands for the Salinas dataset across various target numbers of bands. Each horizontal ”row” represents a selection of bands in a single run of EHBS, with the selected bands marked in black. The top row corresponds to the selection of 8 bands, and the bottom row corresponds to a run with 48 bands selected.
@@
EHBS Waludi Walumi ISSC BS-NETS Conv BS-NETS FC
@@
EHBS Waludi Walumi ISSC BS-NETS Conv BS-NETS FC
@@
EHBS Waludi Walumi ISSC BS-NETS Conv BS-NETS FC
@@
EHBS Waludi Walumi ISSC BS-NETS Conv BS-NETS FC
@@
Figure 7 shows a graph comparing the accuracy of our proposed model, EHBS, as well as the various baseline band selection models on the PaviaU dataset with a patch size of 7x7 over multiple target number of bands. As can be seen from the graph, our proposed model outperforms the baseline models in most of the data points and is comparable to the best baseline model in the others. it also shows the corresponding AUC metric for the various models. Overall, our EHBS model achieved the highest AUC value surpassing the performance of all other models. Figure 8 shows the accuracy results of our proposed method over the whole range of target band selection for the PaviaU dataset and a 7x7 patch size. For this specific dataset, close to optimal performance is obtained by using the 20 best-selected bands (or $\sim$20of the bands). Performance is stable with only slight improvement from this point till the use of the full hyperspectral input.
Table 1 shows a comparison table including accuracy results and AUC results over the PaviaU dataset over two different patch sizes (7x7 and 11x11) and selected target number of bands. As baseline models, we chose the top-performing models from the PaviaU 7x7 experiments. Table 2 shows a similar comparison table for the Salinas dataset. For all datasets and patch sizes, EHBS obtained the highest AUC score and the best accuracy for all Salinas experiments and for the vast majority of PaviaU experiments. In the few PaviaU experiments for which EHBS was not the top-performing method, it obtained comparable accuracy results to the top-performing method.
Figure 6 illustrates the selected bands for the Salinas dataset across various target numbers of bands. EHBS systematically ignores areas with low contribution for the downstream task performance, and the prominent band selection areas remain consistent across different runs.
@@ 6 Discussion and Conclusion
Following, we highlight several key findings and implications derived from our experiments. Firstly, our proposed method demonstrates clear superiority across two diverse datasets and various experimental settings, underscoring its robustness and effectiveness. We conducted comparisons among the methods in various settings, including different patch sizes, datasets, and the number of selected bands. We observed that for larger patch sizes, our method outperformed others by an even wider margin. We attribute this phenomenon to the presence of more contextual information and higher model complexity. Consequently, employing an embedded model enables the model to discover the optimal bands that effectively capture the intricate non-linear interactions among the features. Additionally, our findings suggest that our model tends to outperform baseline models when confronted with an abundance of training data and a higher number of bands to select. We thus hypothesize that our method would perform even better on larger datasets, with larger patch sizes, and with different architectures that hold a global view per band. However, exploration of these scenarios will be reserved for future work. We demonstrated the strength of our model while integrating it into a CNN-based deep learning model for hyperspectral semantic segmentation. The seamless integration of our method into an existing deep learning model that was achieved by simply adding a selection layer after the input enhances its practicality and ease of use. This simplicity of implementation, coupled with the robust performance, positions our method as a promising and practical solution for various machine learning and computer vision tasks.
@@
These results also demonstrate the robustness of our proposed method along multiple settings and various target numbers of bands. The method performs well and consistently without the need to tweak or modify the model regardless of the target number of bands—be it only a few or when aiming for a large number of selected bands. This also allows researchers to get a good understanding of the performance curve and choose the desired tradeoff given the application setting.
@@
After our initial experimentation, we further tested more optimizers by applying an extensive search for an optimal learning rate, as can be seen in Table 3. When using the Adam optimizer, the results are sensitive to the learning rate, and finding the right learning rate is crucial. It turns out that on the PaviaU dataset with a 7x7 path, EHBS with the Adam optimizer can achieve even better performance than EHBS with DoG. However, this is a retrospective result obtained over our test set and not via tuning on a held-out evaluation set. Given that the results achieved with the DoG optimizer are comparable to the optimal outcomes obtained using the Adam optimizer, we advocate for the adoption of DoG. Notably, DoG eliminates the need for parameter tuning, ensuring a more robust and reliable performance.
@@ 7 Acknowledgement
This work was partially supported by the Chief Scientist of the Israeli Ministry of Agriculture grant number 12-03-0010. We would also like to thank Nadav Spitzer for his help in implementing the baseline feature selection models and Ofir Lindenbaum and Jacob Goldberger for providing input and feedback on our work.